OpenStudy (anonymous):

ok question. this one is tricky. my dad and i cant quite figure this out too. please help. i will give medals Maneesh paid $63 in interest to borrow $3600 for 6 months. What annual interest rate did Maneesh pay? A.3.5% B.1.75% C.0.035% D.0.0175%

OpenStudy (anonymous):

please help

OpenStudy (whpalmer4):

Simple interest, or compound?

OpenStudy (whpalmer4):

Do you understand my question? Do you know the difference between simple interest and compound interest?

OpenStudy (whpalmer4):

It appears from working the problem that they are talking about simple interest here, not compound.

OpenStudy (whpalmer4):

Why do you keep asking other people for help when there is someone already helping you?

OpenStudy (anonymous):

nvm tht it is simple intrest

OpenStudy (whpalmer4):

Are you interested in my help, or not?

OpenStudy (the_fizicx99):

Whpalmer is better at explaining than I am :P

OpenStudy (anonymous):

what do u think. of course i am it is just i just burnt myself really bad and i m not in the best mood. @whpalmer4 continue

OpenStudy (anonymous):

oh and i burnt myself on the dishwasher coils

OpenStudy (whpalmer4):

Okay, the formula for computing simple interest is \[I = Prt\] where \(I\) is the total interest paid, \(P\) is the principal balance, \(r\) is the interest rate, expressed as a decimal, and \(t\) is the number of years

OpenStudy (whpalmer4):

Sorry to hear about the burn! No, @shadow22, that's not correct.

OpenStudy (anonymous):

so multiply the three?

OpenStudy (whpalmer4):

Here we have \(I = 63\), \(P=3600\). One tricky bit: what do we use for \(t\)?

OpenStudy (anonymous):

6*2= t

OpenStudy (anonymous):

6 months

OpenStudy (anonymous):

T is in years

OpenStudy (anonymous):

oh ya..... so it would be 0.5?

OpenStudy (anonymous):

0.5=t?

OpenStudy (whpalmer4):

yes, \(t=0.5\) is the correct value to use, because 6 months is 0.5 years

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

now, can you rearrange our equation to give \(r\) in terms of the other three quantities?

OpenStudy (anonymous):

yes... OS is lagging so much don't let me type

OpenStudy (anonymous):

\[I =63 and P=3600 and T=0.5sowhatdoesRmean\]

OpenStudy (anonymous):

and what do i do next

OpenStudy (whpalmer4):

Can you solve \[I = Prt\] for \(r\) on one side and everything else on the other?

OpenStudy (anonymous):

r=rate

OpenStudy (anonymous):

what do you meanr on one side and everything else on the other?

OpenStudy (anonymous):

mean r*

OpenStudy (anonymous):

interest rate

OpenStudy (whpalmer4):

on one side of the = sign

OpenStudy (anonymous):

Interest: $126 for 1 year. Rate: x% of 3,600 = 126 x = 126/3,600 = 0.035 -> the rate is 3.5%

OpenStudy (whpalmer4):

\[I = Prt\]\[\frac{I}{P} = rt\]\[\frac{I}{Pt} = r\]

OpenStudy (anonymous):

i dont get what you mean whpalmer

OpenStudy (whpalmer4):

all three of those formulas are equivalent. In the second one, I divided both sides by \(P\). In the third one, I divided both sides by \(P\) and also by \(t\). The third one is the one I was hoping you could find for me. If you plug in the numbers in the third equation, you'll get your answer, though remember that the value of \(r\) you get is a decimal, NOT a percentage. You'll have to multiply it by 100% to get a percentage.

OpenStudy (anonymous):

i dont get what you mean by "sides"

OpenStudy (anonymous):

:/

OpenStudy (whpalmer4):

sides of the equation. In the initial equation, you have \[I = Prt\]so \(I\) is on the left side and \(Prt\) is on the right side However, to find the value of \(r\), we need \(r\) alone on one side of the = sign, and everything else on the other side. \[I = Prt\]divide both sides by \(P\) \[\frac{I}{P} = \frac{Prt}{P}\]\[\frac{I}{P} = \frac{\cancel{P}rt}{\cancel{P}} = rt\] Any question about that?

OpenStudy (anonymous):

ya i still am confused. i am in a sort of young grade (6) so i dont really understad that equation you made for an example

OpenStudy (anonymous):

sort of. i only understand a little bit of it

OpenStudy (anonymous):

The answer is c hun

OpenStudy (anonymous):

i still want to learn how to do it and whpalmer is helping me alot

OpenStudy (anonymous):

Oh Ok Thats Good

OpenStudy (anonymous):

continue palmer

OpenStudy (whpalmer4):

Well, let's try it without rearranging the formula. \[I = Prt\]We know \(I=63, \,P=3600,\,t=0.5\) That means we have \[63 = 3600*r*0.5\]We can simplify that to \[63 = 1800*r\] If I tell you that 1800 * some number gives you 63, how would you find out what that number is?

OpenStudy (anonymous):

i would divide 1800 by 63

OpenStudy (anonymous):

would i be correct @whpalmer4 ?

OpenStudy (anonymous):

umm....... did everyone just turn dead? or am i lagging

OpenStudy (anonymous):

i have 5 more minutes left

OpenStudy (anonymous):

would i be correct if i said c?

OpenStudy (anonymous):

i am just going with it. good bye

OpenStudy (anonymous):

Oh Sorry Hun Yes Its c

OpenStudy (whpalmer4):

sorry, OpenStudy not being cooperative 63/1800 is the calculation you would make to find the number which when multiplied by 1800 gives you 63. 63/1800 = 0.035 0.035 * 100% = 3.5% To check the answer, 3.5% annual interest means for every $100 borrowed for 1 year, you pay $3.50 in interest. If you only borrow it for half that time (6 months), then you pay half that amount of interest, or $1.75. You borrowed $3600, which can be thought of as borrowing $100 36 times. For each $100, you pay $1.75 for half a year's interest. $1.75 * 36 = $63, which is the total amount the guy paid in interest, so 3.5% is the correct interest rate.

OpenStudy (the_fizicx99):

Yes, you would divide both sides by 1800, \(\ \dfrac{63}{1800} = \dfrac{\cancel{1800}}{\cancel{1800}} \)

OpenStudy (the_fizicx99):

My latex is so laggy :/

OpenStudy (whpalmer4):

C is NOT correct If C was the correct answer, the amount of interest paid for borrowing $3600 for half a year would be only $0.63, not $63.

OpenStudy (the_fizicx99):

@Michelle♥ forget to convert it back by multiplying by 100 >.>

OpenStudy (whpalmer4):

Thus illustrating a danger of accepting answers from people who don't show their work!

OpenStudy (whpalmer4):

Anyhow, if you understand a bit of algebra, you would rearrange the formula I gave: \[I = Prt\]by dividing both sides by \(Pt\) to give\[\frac{I}{Pt} = r\]Plugging in the known values \[\frac{\$63}{(\$3600)(0.5)} = 0.035\]Again, this is a decimal rate, not a percentage, so we multiply by \(100\%\) to get \(3.5\%\) as the annual interest rate for the loan.